a regular polygon has 14 sides. what is the measure of each vertex?

Answer by Alan3354(69443)

a regular polygon has 14 sides. what is the measure of each vertex?
-------------
The sum of external angles is always 360 degs
Each ext angle = 360/14 = 180/7 degs
Internal angles = 180 - 180/7 = 1080/7 degs
=~ 154.286 degs